OpenStudy (anonymous):

need help on this problem i got the answer 10^(1/2) and -10^(1/2) but im not sure.. Find the real solution(s) of the equation involving rational exponents. Check your solutions (x^2 + 17)^2/3 = 9

OpenStudy (amistre64):

\[{(x^2 + 17)^2\over3} = 9 \] this equation?

OpenStudy (amistre64):

\[(x^2 + 17)^{2/3} = 9 \]or this one?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

9^(3/2)= 27 so 27= x^2+17 10=x^2 thats what i got

OpenStudy (amistre64):

cube both sides to begin with right? (x^2 +17)^2 = 729 ; sqrt the sides now x^2 +17 = +- sqrt(729) ; -17 x^2 = -17 +- sqrt(729) is what im looking at

OpenStudy (anonymous):

is there a way of writing that as a fraction because my assignment is marking it wrong when i submit it as 10^(1/2)

OpenStudy (amistre64):

sqrt the sides again to get: x = \(\large \pm\sqrt{-17 \pm \sqrt{729}}\) :)

OpenStudy (anonymous):

sqrt of 10

OpenStudy (amistre64):

the only 'real' solution that comes out of the is sqrt(10) = 10^(1/2)

OpenStudy (anonymous):

?

OpenStudy (amistre64):

well, +- sqrt(10); which you had already gotten :)

OpenStudy (amistre64):

what are your options for inputing the answer? is there a sqrt button?

OpenStudy (anonymous):

\sqrt{blah} will give it to you